OpenStudy (anonymous):

Find an equation of the line. Write the equation using function notation. Through (-6,-9); parallel to 5x+6y=11 f(x)=

OpenStudy (anonymous):

@johnweldon1993 Please :(

OpenStudy (johnweldon1993):

Alright first things first...lets get the line you have into slope-intercept form \[\large y = mx + b\] that one lol where 'm' is the slope...and 'b' is the y-intercept 5x + 6y = 11 subtract 5x from both sides 6y = -5x + 11 And divide everything by 6 \[\large y = -\frac{5}{6}x + \frac{11}{6}\] This is the equation of your other line So...what do we know about parallel lines in terms of slopes?

OpenStudy (anonymous):

They are quite complicated... And 2 fractions :)

OpenStudy (johnweldon1993):

Lol :P Alright...well...we know the the slopes of parallel lines...are the same correct? If they were not...they would cross and not be parallel So If the slope of the line we have....is -5/6 then the slope of the line we want...is also -5/6 Make sense so far?

OpenStudy (anonymous):

yes so the first fraction is -5/6 ?

OpenStudy (johnweldon1993):

Correct.... okay now..we need to introduce another form...called the point slope form why? well it will tell us the equation of a line...if we have a point...and a slope :P lol which we do...we are given a point and we just found the slope so it looks like \[\large y - y_0 = m(x - x_0)\] same thing here...m = slope....so it will be the -5/6 we had

OpenStudy (anonymous):

gotcha

OpenStudy (johnweldon1993):

and the \(\large (x_0 , y_0)\) will be replaced by your point \(\large (-6,-9)\) so altogether we will have \[\large y - (-9) = -\frac{5}{6}(x - (-6))\] still with me there?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

So lets simplify it a bit, distribute the -5/6 into the parenthesis \[\large y + 9 = -\frac{5}{6}x - 5\] Then just subtract 9 from both sides \[\large y = -\frac{5}{6}x - 14\] and there we have our equation :)

OpenStudy (anonymous):

thank you so much i really appreciate it!

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

do you wanna help me with one more or nah?

OpenStudy (johnweldon1993):

nah ;P lol no jk go ahead :)